find the slope and the y intercept. Y=4.8x-8. The slope is?
Answer by ewatrrr(24785)


Hi


Y = 4.8x - 8 |slope m = 4.8 OR 48/10 = 24/5 and y-intercept is -8
Line crossing the y axis at Pt(0,-8, slanting to right as slope >0
m = 24/5: as y increases/decreases by 24,
x increases/decreases by 5